OpenStudy (anonymous):

What is the 6th term in this geometric sequence? 480, 240, 120, 60, ...

OpenStudy (anonymous):

480/2 = 240, 240/2 = 120, 120/2 = 60, 60/2 = 30, 30/2 = 15. 15 is the 6th term